# Xu Duan, Lei Sun [^2] Corresponding author. , and Xiangxiang Chu Amap, Alibaba Group

### v2

@@ Abstract
Controllable image synthesis, which enables fine-grained control over generated outputs, has emerged as a key focus in visual generative modeling. However, controllable generation remains challenging for Visual Autoregressive (VAR) models due to their hierarchical, next-scale prediction style. Existing VAR-based methods often suffer from inefficient control encoding and disruptive injection mechanisms that compromise both fidelity and efficiency. In this work, we present SCALAR, a controllable generation method based on VAR, incorporating a novel Scale-wise Conditional Decoding mechanism. SCALAR leverages a pretrained image encoder to extract semantic control signal encodings, which are projected into scale-specific representations and injected into the corresponding layers of the VAR backbone. This design provides persistent and structurally aligned guidance throughout the generation process. Building on SCALAR, we develop SCALAR-Uni, a unified extension that aligns multiple control modalities into a shared latent space, supporting flexible multi-conditional guidance in a single model. Extensive experiments show that SCALAR achieves superior generation quality and control precision across various tasks. All the code will be released.
@@ 1 Introduction
Controllable image synthesis is a pivotal domain in visual generation, enabling the precise and nuanced creation of visual content according to specific user guidance. Recent advances in this field are currently dominated by two primary paradigms: Diffusion Models [11, 41] and Autoregressive Models (AR) [43, 45]. While diffusion-based methods [57, 17, 32] have achieved widespread success, the iterative denoising process is not inherently compatible with the sequential, token-based architecture of LLMs, which hinders the development of truly unified multimodal modeling [61, 51].
@@
Visual Autoregressive (VAR) models [45], a leading approach within the autoregressive paradigm, offer a compelling path forward. By framing image synthesis as a next-scale prediction task, VAR-based models [9, 44, 24, 63] align naturally with LLM architectures and have demonstrated better inference efficiency and generative quality than both state-of-the-art diffusion models [31] and raster-scan-based autoregressive models [43]. Nevertheless, the capacity of VAR models for fine-grained control remains a significant and underexplored challenge. This challenge primarily stems from the unique hierarchical, scale-wise generation process, which presents a distinct paradigm from existing approaches: In diffusion models, control signals are applied globally to the denoising network at each step, whereas in traditional raster-scan AR models [43, 16], they are injected before each spatial token prediction.
@@
Existing works on controllable Visual Autoregressive learning, notably CAR [53] and ControlVAR [20], deliver suboptimal performance. We attribute this bottleneck to two fundamental design flaws. First, they utilize complex and disruptive injection mechanisms; architectures with parallel branches [53] similar to ControlNet or joint control and image modeling schemes [20] not only incur substantial computational overhead but also implicitly disrupt the powerful generative capabilities of the pretrained backbone. Second, they often utilize lightweight convolutional networks or VQ-VAEs as control encoders, which have been proven to have a limited capacity in capturing rich spatial-semantic features [62, 19]. Consequently, their performance in both generation quality and control consistency trails behind even raster-scan AR models [21], highlighting a critical need for a control mechanism that is both efficient and fundamentally aligned with the native structure of VAR.
@@
To this end, we present SCALAR, an effective controllable generation method based on VAR. Aligning with the more efficient paradigm [21], SCALAR injects control during the autoregressive decoding phase. Our approach is centered on a novel and simple Scale-wise Conditional Decoding mechanism. To be specific, we first employ a pretrained vision foundation model [29, 14, 6] to extract powerful, scale-agnostic control features that contain rich semantic information. These general features are then processed by scale-wise lightweight projection blocks—which maintain independent weights for each scale—to produce specialized Control Signal Encodings. Ultimately, SCALAR injects these tailored encodings directly into the hidden states of the VAR backbone’s layers, providing scale-wise persistent guidance throughout the next-scale prediction process. Building on our design, we extend SCALAR to a unified control version, SCALR-uni. This extension incorporates a Unified Control Alignment process that projects features from diverse control modalities into a common latent space, enabling seamless guidance of various condition types from a unified model.
@@
  • Extensive experiments on ImageNet demonstrate that SCALAR and SCALAR-Uni achieve exceptional generation quality and conditional consistency across all five conditional generation tasks (e.g., FID on Canny: 2.14 vs. 7.85; Depth: 3.09 vs. 4.19).
@@ 2.1 Image Generation
Two dominant paradigms have emerged in image generation: diffusion models and AR models. Diffusion models synthesize images by iteratively denoising Gaussian noise, with DDPM [12] marking a major breakthrough. Subsequent work improves generation quality and efficiency [28, 36, 59] through advances in sampling strategies and latent-space modeling. They have become the backbone of text-to-image and text-to-video generation [37, 39, 23, 15, 60], typically using U-Net for denoising and CLIP or T5 for text conditioning via cross-attention. Recent models like DiT [30] replace U-Net with Transformer backbones, achieving strong performance. Despite their success, diffusion models remain computationally expensive, motivating exploration of more efficient alternatives such as autoregressive approaches.
@@
In contrast, AR models formulate image synthesis as a sequence modeling task, predicting tokens step by step. Early works [48] focused on pixel-level generation. Building on the success of large language models [47, 1], emerging approaches adopt discrete quantizers such as VQ-VAE [49] and VQ-GAN [7] to convert image patches into discrete token indices, enabling next-token prediction over visual token sequences. Representative works such as LlamaGen [43] and Open-MAGVIT2 [22] adopt the LLaMA architecture [46], while AiM [16] explores the Mamba [8] as the autoregressive backbone. Among them, Visual Autoregressive Modeling (VAR) introduces a scalable next-scale prediction mechanism, which differs from traditional raster-scan AR methods. VAR-based models for class-to-image and text-to-image tasks [45, 24, 44, 9, 63] achieve image synthesis performance comparable to state-of-the-art diffusion models while offering significant computational efficiency.
@@ 2.2 Controllable Image Generation
Conditional image generation, where the model is guided by external signals during the generative process, has become an active area of research. While early efforts leveraged class labels or attributes in GANs [25, 40] and VAEs [13, 56], recent advances focus on more fine-grained and flexible conditioning. In diffusion models, representative frameworks [57, 27] incorporate these signals via cross-attention or adapter-based modules, enabling high-quality, controllable generation without the need to fine-tune the full model. Follow-up works [32, 17, 34] further improve generalization and adaptability across diverse conditions and architectures.
In contrast, conditional generation for AR models has been far less explored. Existing works on controllable autoregressive learning are typically classified according to the guidance injection phase in autoregressive generation: pre-filling phase or decoding phase. Pre-filling-phase methods [20, 33] fill the control encodings into the initial sequence from the start. This approach increases the token sequence length, inflating the computational load, while the joint modeling scheme in ControlVAR [20] can disrupt the powerful capabilities of the pretrained backbone. Decoding-phase methods [21], which continuously inject control encodings throughout the step-by-step decoding, are simpler and more flexible. However, CAR [53] adopts parallel conditional decoding branches inspired by ControlNet, which introduces huge complexities when autoregressive decoding. Notably, although VAR models generally outperform traditional AR models in image synthesis quality, these VAR-based controllable methods [20, 53] still underperform compared to raster-scan AR-based methods [21] in terms of control consistency and image quality. This gap indicates that the potential of VAR in controllable generation is yet to be fully explored. In this paper, we aim to develop a general and efficient method for controllable image generation based on VAR.
@@ 3 SCALAR
In this section, we present SCALAR, a simple and efficient method for controllable VAR. We first revisit the image generation with VAR. Next, we introduce Scale-wise Conditional Decoding. Then, we conduct several experiments to find the model architecture that yields the best generation quality and control consistency. Finally, we extend our approach to a unified version, SCALR-Uni, with the proposed Unified Control Alignment.
Figure 3: (a) Comparison of parameter sharing for projection blocks ($\mathcal{P}_{k,l}$, $\mathcal{P}_{k}$, and $\mathcal{P}_{l}$). (b) Comparison of various injection layers set ($\mathcal{S}_{1}$, $\mathcal{S}_{\text{alt}}$, and $\mathcal{S}_{\text{all}}$) with different structures of projection block (Linear and LinearLite). (c) Comparison of different parameter-efficient training strategies ($\text{Frz}_{\text{none}}$, $\text{Frz}_{\text{SA}}$, and $\text{Frz}_{\text{all}}$). (d) Impacts of scaling up the depth of VAR backbone (VAR-d12, d16, and d20).
@@ 3.1 Preliminary
Building upon the success of large language models [47, 1], visual autoregressive models adopt discrete quantizers such as VQVAE to convert image patches into index-wise tokens, enabling image generation via next-step prediction over visual token sequences. Among them, VAR introduces a scalable next-scale prediction mechanism, which generates images in a coarse-to-fine manner across spatial resolutions. VAR-based class-to-image (c2i) [45] and text-to-image (t2i) [24, 44, 9, 63] models achieve image synthesis performance comparable to state-of-the-art diffusion models, while reducing computational cost.
The core idea of the VAR model lies in next-scale prediction, which contrasts with traditional raster-scan AR models based on next-token prediction. While conventional AR models generate images token by token along a flattened pixel sequence, this formulation suffers from mathematical inconsistencies and often leads to structural degradation, especially in highly structured images. In contrast, VAR avoids these issues by operating at the level of token maps, predicting the image progressively across multiple scales. Starting from a coarse $\times 1$ 1 token map $r_{1}$, it autoregressively generates a sequence of higher-resolution token maps $(r_{2},\dots,r_{K})$, each representing a finer level of detail. The overall generation process is formulated as:
@@ 3.2 Scale-wise Conditional Decoding
The generation process in autoregressive models is bifurcated into two phases: pre-filling and decoding. In SCALAR, we diverge from the conditional pre-filling way [20, 33], which relies on a distinct initial conditioning filling step. We adopt a simpler conditional decoding strategy [21, 53]. This design integrates the control signal directly and continuously throughout the decoding process. The scale-autoregressive nature of VARs poses distinct requirements for how control signals are applied across the generative hierarchy; guidance must be present and adapted at each scale. SCALAR achieves this by injecting the control signal into a predefined subset of layers, indexed by $l\in\mathcal{S}$. Formally, the operation at each layer is expressed as:
@@
Control Signal Encoding $\mathbf{C}_{k,l}$. We employ a pretrained vision foundation model [29, 14, 6] as our universal control signal encoder. Its inherent understanding of rich visual semantics, learned via large-scale self-supervision, provides a powerful and robust feature extractor for diverse control conditions. To leverage the hierarchical features within the encoder [2], the control representation $\mathbf{F}_{c}$ is formed by concatenating features from the $i$-th layers of the encoder $\mathcal{E}$. It can be expressed as:
@@
where $\mathcal{E}_{i}(\cdot)$ is the feature from the $i$-th layer of the encoder, $\mathcal{C}(\cdot)$ represents the concatenation operation and $\mathcal{I}$ is the set of indices for the selected layers. The resulting scale-agnostic feature $\mathbf{F}_{c}$ is then tailored for injection into scale $k$ and layer $l$ of our Scale-wise Conditional Decoding process
@@
Table 1: Quantitative results of conditional image generation on ImageNet [5]. Values marked with ~ are estimated from histograms in the corresponding paper. Cells highlighted with xx indicate the best performance for each setting.
@@
Diff. T2IAdapter [26] - ~10.2 ~157 - ~9.9 ~134 - ~9.5 ~143 - ~9.3 ~142 - ~16.2 ~156 -
ControlNet [58] - ~11.6 ~173 - ~9.2 ~150 - ~8.9 ~155 - ~8.6 ~150 - ~15.3 ~163 -
AR ControlAR [21] AiM-L 9.66 - 30.36 7.39 - 35.01 - - -
@@ 3.3 Exploring Controlled Architectural Design
[…]

### v3

@@ Abstract
Controllable image synthesis, which enables fine-grained control over generated outputs, has emerged as a key focus in visual generative modeling. However, controllable generation remains challenging for Visual Autoregressive (VAR) models due to their hierarchical, next-scale prediction style. Existing VAR-based methods often suffer from inefficient control encoding and disruptive injection mechanisms that compromise both fidelity and efficiency. In this work, we present SCALAR, a controllable generation method based on VAR, incorporating a novel Scale-wise Conditional Decoding mechanism. SCALAR leverages a pretrained image encoder to extract semantic control signal encodings, which are projected into scale-specific representations and injected into the corresponding layers of the VAR backbone. This design provides persistent and structurally aligned guidance throughout the generation process. Building on SCALAR, we develop SCALAR-Uni, a unified extension that aligns multiple control modalities into a shared latent space, supporting flexible multi-conditional guidance in a single model. quality and control precision, including diffusion-based methods (e.g., ControlNet++) and AR-based methods (e.g., ControlAR, CAR). The code is released at https://github.com/AMAP-ML/SCALAR.
@@ 1 Introduction
Controllable image synthesis is a pivotal domain in visual generation, enabling the precise and nuanced creation of visual content according to specific user guidance. Recent advances in this field are currently dominated by two primary paradigms: Diffusion Models [ddpm, song2020score] and Autoregressive Models (AR) [llamagen, var]. While diffusion-based methods [controlnet, controlnet++, unicontrol] have achieved widespread success, the iterative denoising process is not inherently compatible with the sequential, token-based architecture of LLMs, which hinders the development of truly unified multimodal modeling [zhao2023easygen, wang2024multi].
@@
Visual Autoregressive (VAR) models [var], a leading approach within the autoregressive paradigm, offer a compelling path forward. By framing image synthesis as a next-scale prediction task, VAR-based models [han2025infinity, tang2024hart, ma2024star, zhuang2025vargpt] align naturally with LLM architectures and have demonstrated better inference efficiency and generative quality than both state-of-the-art diffusion models [sdxl] and raster-scan-based autoregressive models [llamagen]. Nevertheless, the capacity of VAR models for fine-grained control remains a significant and underexplored challenge. This challenge primarily stems from the unique hierarchical, scale-wise generation process, which presents a distinct paradigm from existing approaches: In diffusion models, control signals are applied globally to the denoising network at each step, whereas in traditional raster-scan AR models [llamagen, aim], they are injected before each spatial token prediction.
@@
Existing works on controllable Visual Autoregressive learning, notably CAR [yao2024car] and ControlVAR [controlvar], deliver suboptimal performance. We attribute this bottleneck to two fundamental design flaws. First, they utilize complex and disruptive injection mechanisms; architectures with parallel branches [yao2024car] similar to ControlNet or joint control and image modeling schemes [controlvar] not only incur substantial computational overhead but also implicitly disrupt the powerful generative capabilities of the pretrained backbone. Second, they often utilize lightweight convolutional networks or VQ-VAEs as control encoders, which have been proven to have a limited capacity in capturing rich spatial-semantic features [zhou2025dino, li2024imagefolder]. Consequently, their performance in both generation quality and control consistency trails behind even raster-scan AR models [li2024controlar], highlighting a critical need for a control mechanism that is both efficient and fundamentally aligned with the native structure of VAR.
@@
To this end, we present SCALAR, an effective controllable generation method based on VAR. Aligning with the more efficient paradigm [li2024controlar], SCALAR injects control during the autoregressive decoding phase. Our approach is centered on a novel and simple Scale-wise Conditional Decoding mechanism. To be specific, we first employ a pretrained vision foundation model [dinov2, sam, vit] to extract powerful, scale-agnostic control features that contain rich semantic information. These general features are then processed by scale-wise lightweight projection blocks—which maintain independent weights for each scale—to produce specialized Control Signal Encodings. Ultimately, SCALAR injects these tailored encodings directly into the hidden states of the VAR backbone’s layers, providing scale-wise persistent guidance throughout the next-scale prediction process. Building on our design, we extend SCALAR to a unified control version, SCALAR-uni. This extension incorporates a Unified Control Alignment process that projects features from diverse control modalities into a common latent space, enabling seamless guidance of various condition types from a unified model.
@@
  • Extensive experiments demonstrate that our proposed SCALAR exhibits exceptional generation quality and conditional consistency across both class-to-image (c2i) and text-to-image (t2i) controllable generation, significantly outperforming other SoTA methods.
@@ 2.1 Image Generation
Among recent advances in image generation, two paradigms have gained particular traction: diffusion models and autoregressive (AR) models. Diffusion-based methods, beginning with DDPM [ho2020denoising], approach generation by progressively denoising Gaussian noise. This framework has since evolved significantly, with improvements in sampling techniques and latent representations [nichol2021improved, rombach2022high, zhang2025robust, peng2024efficient, peng2025pixel] boosting both visual quality and efficiency. Such models now power the majority of high-fidelity text-to-image and text-to-video systems [saharia2022photorealistic, singer2022make, ma2024passersby, lan2025flux, peng2025boosting, zhang2025boow, ragsr, peng2024unveiling], typically adopting a U-Net architecture for denoising and using CLIP or T5 as the backbone for text understanding. More recently, diffusion transformers like DiT [dit] have demonstrated the scalability of replacing U-Nets with Transformer backbones. Despite their success, diffusion models are known for high computational costs, prompting a shift of attention toward AR models as a more efficient alternative.
@@
Autoregressive (AR) models take a different route, casting image generation as a sequential prediction task. Initial methods [van2016pixel] directly modeled raw pixel outputs in raster-scan order. Recent developments, inspired by the progress of large language models [touvron2023llama, achiam2023gpt], introduce discrete visual representations using tokenizers like VQ-VAE [van2017neural] and VQ-GAN [esser2021taming], allowing AR models to operate on compact image tokens and predict them autoregressively. Recent works [an2025unictokens, an2024mc, luo2024llm, lin2025perceive, lin2024draw] such as LlamaGen [llamagen] and Open-MAGVIT2 [open-magvit2] adopt the LLaMA architecture [llama], while AiM [aim] explores the Mamba [mamba] as the autoregressive backbone. Among them, Visual Autoregressive Modeling (VAR) introduces a scalable next-scale prediction mechanism, which differs from traditional raster-scan AR methods. VAR-based models for class-to-image and text-to-image tasks [var, ma2024star, tang2024hart, han2025infinity, zhuang2025vargpt] achieve image synthesis performance comparable to state-of-the-art diffusion models while offering significant computational efficiency.
@@ 2.2 Controllable Image Generation
Conditional image generation, where models are guided by external signals, has become a central topic in generative modeling. Early efforts focused on class labels and attributes, explored via GANs [mirza2014conditional, song2023fashion] and VAEs [kingma2013auto, zeng2024cat]. More recent diffusion-based methods [controlnet, t2i-adapter] integrate control signals through cross-attention or adapter modules, enabling high-quality generation without full-model fine-tuning. Follow-up work [unicontrol, controlnet++, ran2024x] further improves generalization across control types and model architectures.
In contrast, conditional generation for AR models has been far less explored. Existing works on controllable autoregressive learning are typically classified according to the guidance injection phase in autoregressive generation: pre-filling phase or decoding phase. Pre-filling-phase methods [controlvar, qu2025varsr] fill the control encodings into the initial sequence from the start. This approach increases the token sequence length, inflating the computational load, while the joint modeling scheme in ControlVAR [controlvar] can disrupt the powerful capabilities of the pretrained backbone. Decoding-phase methods [li2024controlar], which continuously inject control encodings throughout the step-by-step decoding, are simpler and more flexible. However, CAR [yao2024car] adopts parallel conditional decoding branches inspired by ControlNet, which introduces huge complexities when autoregressive decoding. Notably, although VAR models generally outperform traditional AR models in image synthesis quality, these VAR-based controllable methods [controlvar, yao2024car] still underperform compared to raster-scan AR-based methods [li2024controlar] in terms of control consistency and image quality. This gap indicates that the potential of VAR in controllable generation is yet to be fully explored. In this paper, we aim to develop a general and efficient method for controllable image generation based on VAR.
@@ 3 SCALAR
In this section, we present SCALAR, a simple and efficient method for controllable VAR. We first revisit the image generation with VAR. Next, we introduce Scale-wise Conditional Decoding. Then, we conduct several experiments to find the model architecture that yields the best generation quality and control consistency. Finally, we extend our approach to a unified version, SCALAR-Uni, with the proposed Unified Control Alignment.
Figure 3: (a) Comparison of parameter sharing for projection blocks ($\mathcal{P}_{k,l}$, $\mathcal{P}_{k}$, and $\mathcal{P}_{l}$). (b) Comparison of various injection layers set ($\mathcal{S}_{1}$, $\mathcal{S}_{\text{alt}}$, and $\mathcal{S}_{\text{all}}$) with different structures of projection block (Linear and LinearLite). (c) Comparison of different parameter-efficient training strategies ($\text{Frz}_{\text{none}}$, $\text{Frz}_{\text{SA}}$, and $\text{Frz}_{\text{all}}$). (d) Impacts of scaling up the depth of VAR backbone (VAR-d12, d16, and d20). Note: All experiments are conducted on ImageNet [imagenet] with the c2i controllable generation.
@@ 3.1 Preliminary
Building upon the success of large language models [touvron2023llama, achiam2023gpt], autoregressive models adopt discrete quantizers such as VQVAE to convert image patches into index-wise tokens, enabling image generation via next-token prediction over visual token sequences. Among them, VAR introduces a scalable next-scale prediction mechanism to generate images across resolutions. VAR-based class-to-image (c2i) [var] and text-to-image (t2i) [ma2024star, tang2024hart, han2025infinity, zhuang2025vargpt] models achieve image synthesis performance comparable to state-of-the-art diffusion models, while reducing computational cost.
The core idea of the VAR model lies in next-scale prediction, which contrasts with traditional raster-scan AR models based on next-token prediction. While conventional AR models generate images token by token along a flattened pixel sequence, this formulation suffers from mathematical inconsistencies and often leads to structural degradation, especially in highly structured images. In contrast, VAR avoids these issues by operating at the level of token maps, predicting the image progressively across multiple scales. Starting from a coarse $1\times 1$ token map $r_{1}$, it autoregressively generates a sequence of higher-resolution token maps $(r_{2},\dots,r_{K})$, each representing a finer level of detail. The overall generation process is formulated as:
@@ 3.2 Scale-wise Conditional Decoding
The generation process in autoregressive models is bifurcated into two phases: pre-filling and decoding. In SCALAR, we diverge from the conditional pre-filling way [controlvar, qu2025varsr], which relies on a distinct initial conditioning filling step. We adopt a simpler conditional decoding strategy [li2024controlar, yao2024car]. This design integrates the control signal directly and continuously throughout the decoding process. The scale-autoregressive nature of VARs poses distinct requirements for how control signals are applied across the generative hierarchy; guidance must be present and adapted at each scale. SCALAR achieves this by injecting the control signal into a predefined subset of layers, indexed by $l\in\mathcal{S}$. Formally, the operation at each layer is expressed as:
@@
Control Signal Encoding $\mathbf{C}_{k,l}$. We employ a pretrained vision foundation model [dinov2, sam, vit] as our universal control signal encoder. Its inherent understanding of rich visual semantics, learned via large-scale self-supervision, provides a powerful and robust feature extractor for diverse control conditions. To leverage the hierarchical features within the encoder [bolya2025meta_perception_encoder], the control representation $\mathbf{F}_{c}$ is formed by concatenating features from the $i$-th layers of the encoder $\mathcal{E}$. It can be expressed as:
@@
where $\mathcal{E}_{i}(\cdot)$ is the feature from the $i$-th layer of the encoder, $\mathcal{C}(\cdot)$ represents the concatenation operation and $\mathcal{I}$ is the set of indices for the selected layers. The resulting scale-agnostic feature $\mathbf{F}_{c}$ is then tailored for injection into scale $k$ and layer $l$ of our Scale-wise Conditional Decoding process:
@@
Table 1: Quantitative results of conditional image generation on ImageNet [deng2009imagenet]. Values marked with ~ are estimated from histograms in the corresponding paper. Cells highlighted with $x$ indicate the best performance for each setting.
@@
Diff. T2IAdapter [mou2024t2i] - ~10.2 ~157 - ~9.9 ~134 - ~9.5 ~143 - ~9.3 ~142 - ~16.2 ~156 -
ControlNet [zhang2023adding] - ~11.6 ~173 - ~9.2 ~150 - ~8.9 ~155 - ~8.6 ~150 - ~15.3 ~163 -
AR ControlAR [li2024controlar] AiM-L 9.66 - 30.36 7.39 - 35.01 - - -
@@ 3.3 Exploring Controlled Architectural Design
For all comparisons above, we train VAR-d12 models on ImageNet $256\times 256$ with batch size 512. We train models with the above changes and compare them on Fréchet Inception Distance (FID), Inception Score (IS), Root Mean Square Error (RMSE, conditioned on Depth), and F1-Score (conditioned on Canny). As shown in Figure 3(a) and (b), the optimal configuration employs projection blocks with scale-wise parameters for each block ($\mathcal{P}_{k,l}$), uses a standard linear layer (Linear), and injects the control signal into all layers ($\mathcal{S}_{\text{all}}$). Figure 3(c) shows that freezing backbone parameters causes a significant reduction of control performance. In Figure 3(d), scaling up depth of VAR helps performance. Accordingly, we adopt the best combination of settings ($\mathcal{P}_{k,l}$-Linear-$\mathcal{S}_{\text{all}}$-$\text{Frz}_{\text{none}}$) for our final SCALAR model to achieve the best performance. Further analysis details about Figure 3 are provided in the Supplementary Material.